OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (6, 0) and a directrix of y = −10 ?

OpenStudy (anonymous):

These are the options i dont really understand what to do. f(x) = −one twentieth (x − 6)2 + 5 f(x) = −one twentieth (x − 6)2 − 5 f(x) = one twentieth (x − 6)2 + 5 f(x) = one twentieth (x − 6)2 − 5

OpenStudy (aum):

The graph will be a parabola. Parabola has this property where any point (x,y) on the parabola will be equidistant from the focus and the directrix. Therefore, (x-6)^2 + (y-0)^2 = (y - -10)^2 (x-6)^2 + y^2 = (y +10)^2 simplify

OpenStudy (aum):

(x-6)^2 + y^2 = (y +10)^2 (x-6)^2 = (y +10)^2 - y^2 = (y+10+y)(y+10-y) = (2y+10)(10) = 2(y+5)(10) = 20(y+5) 1/20*(x-6)^2 = y+5 y = 1/20*(x-6)^2 - 5

OpenStudy (anonymous):

thank you!

OpenStudy (aum):

you are welcome.